A box is filled with 5 green cards, 2 yellow cards, and 11 blue cards. A card is chosen at random from the box. What is the probability that it is a green or a blue card?
Answer by stanbon(75887)

A box is filled with 5 green cards, 2 yellow cards, and 11 blue cards. A card is chosen at random from the box.
What is the probability that it is a green or a blue card?
-----
P(green or blue) = (5+11)/18 = 16/18 = 8/9
==================
Cheers,
Stan H.
==================